OpenStudy (anonymous):

find the solution for the equation |2x-8|+2=7

OpenStudy (anonymous):

how do you find the solution set

OpenStudy (mr.math):

\[|2x-8|+2=7 \implies |2x-8|=5.\] This is true for two cases: Case 1: \(2x-8=5 \implies 2x=13 \implies x=\frac{13}{2}\). Case 2: \(2x-8=-5 \implies 2x=3 \implies x=\frac{3}{2}\).

OpenStudy (across):

If you're stuck, you could remember the definition\[|a|=b\implies-b=a=b.\]

OpenStudy (anonymous):

|2x -8|+2=7 so, either (2x-8)+2 = 7 or -(2x-8)+2 = 7 2x-6=7 -2x+10=7 2x=13 2x=3 x=13/2 x=3/2